Calculate the mass of a 3cm cube of aluminium.
I think the question is based on the previous that was in my book. It said: a 2cm cube of aluminium has mass 21.6g. It's density is 2.7g/cm<cubed>
1
Expert's answer
2012-01-20T07:45:25-0500
M = P*V, where P=2.7g/cm&sup3; is the density of an aluminium

If the length of cube edge is 3 cm then volume of such cube is 3*3*3=27 cm3
Now substituting value of volume we get

M = 2.7g/cm&sup3;*27cm&sup3; = 72.9g.